### v3

# Complex Frequency Detection in a Subsystem

Abstract
In this study, we systematically explore the non-Hermitian skin effect (NHSE) and its associated complex-frequency detection in the context of a frequency-dependent non-Hermitian Hamiltonian. This Hamiltonian arises from the self-energy correction of the subsystem and can be calculated exactly within our theoretical model, without the need for any approximations. Additionally, complex frequency detection, which encompasses complex frequency excitation, synthesis, and fingerprint, enables us to detect the physical responses induced by the complex driving frequencies. Our calculations reveal that both complex frequency excitation and synthesis are incompatible with the non-Hermitian approximation and are unable to characterize the presence or absence of the NHSE. In contrast, the complex-frequency fingerprint successfully detects the novel responses induced by the NHSE through the introduction of a double-frequency Green’s function. Our work provides a platform for studying non-Hermitian physics and their novel response in quantum systems rigorously without relying on any approximations.
Introduction.—Non-Hermitian physics, particularly the non-Hermitian skin effect (NHSE), has attracted significant research interest recently [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Many experimental developments have been made in classical wave platforms, including acoustic [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], optical [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66], mechanical [67, 68, 69, 70, 71, 72, 73], and electrical circuit systems [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. For quantum systems, a natural theoretical framework involves starting with the standard Green’s function (GF) of subsystems [32, 87, 88, 89, 90, 91]. As schematized in Fig. 1 (a), when both excitations and responses are confined within a subsystem $S$, a frequency-dependent self-energy emerges, i.e., $\Sigma_{S}(k,\omega)$. Under zeroth-order approximation, i.e., $\Sigma_{S}(k,\omega)\simeq\Sigma_{S,0}$, this yields an effective non-Hermitian Hamiltonian, i.e., $H_{S,{\rm nH}}(k)=H_{S}(k)+\Sigma_{S,0}$, referred to as the non-Hermitian approximation (NHA) in this work [92, 93, 88, 87, 94, 95, 96, 97, 98, 99].
Figure 1: (a) Schematic illustration of the excitation and detection processes within the subsystem $S$, which is coupled to another subsystem $S^{\prime}$. (b) Energy spectra of the approximated non-Hermitian $H_{S,\rm{nH}}$ under PBC and OBC. (c) Comparative analysis of the CFGF obtained with and without the NHA. Detailed model parameters and computational specifications are provided in Fig.2.
When $H_{S,\rm{nH}}(k)$ exhibits non-zero spectral winding as shown in Fig. 1 (b), the corresponding open boundary condition (OBC) system manifests the NHSE [26, 30]. In order to understand this phenomenon, the generalized Brillouin zone or non-Bloch band theories are developed [7, 22, 31, 30]. Physical responses governed by Brillouin/generalized Brillouin zones are categorized as Bloch/non-Bloch responses, respectively [9, 100, 101, 102]. For instance, the single-particle GF under OBC, i.e., $G_{\rm OBC}^{R}(\omega)$, belongs to the non-Bloch responses when $\omega$ possesses non-zero spectral winding and Bloch responses otherwise [103]. Fig. 1 (c) illustrates this statement: frequencies from $\omega_{3}$ to $\omega_{5}$ (with non-zero spectral winding) exhibit rightward-dominating divergence in the GF upon excitation at $x=40$. Such a divergent behavior in the GF is a hallmark of the non-Bloch response and can be used to characterize the presence of point gap topology as well as the NHSE [26, 30, 103].
Figure 2: (a) Schematic representation of the tight-binding model for the composite system, with subsystem $S$ explicitly demarcated by the red rectangular enclosure. (b) and (c) A comparison between the self-energy and the bare bands of the subsystem $S$, which ensures the validity of the NHA. The parameters include $t_{1}=1.2,t_{2}=-1,\lambda=-1,\mu_{S}=0.2,t_{x}=1,t_{y}=3.5,\mu_{S^{\prime}}=0.3,t_{A}=0.6,\gamma=0.1$, $N_{S}=80$ and $N_{y}=300$.
In this work, we address the following crucial question: Does the non-Bloch response persist in the exact GF beyond the NHA framework? Through a concrete model (detailed later), Fig. 1 (c) (right panel) demonstrates that the unidirectional divergence becomes absent in the exact GF without applying any approximations. This result raises three fundamental issues: (i) How to interpret this result? (ii) Does our subsystem truly exhibit the NHSE and non-Bloch responses? (iii) If so, how do we detect them? Notably, since the frequency $\omega$ in the GF has become complex, it is essential to explore complex frequency detection methods to answer these questions.
Complex frequency detection techniques, comprising the complex frequency excitation (CFE) [104, 105, 106, 107, 108, 109, 110, 111, 112, 113], synthesis (CFS) [104, 105, 114], and fingerprinting (CFF) [100], offer experimentally viable strategies to extract critical information from the complex-frequency GF (CFGF) [103]. However, existing studies typically assume a frequency-independent non-Hermitian Hamiltonian, i.e., $H_{S,\rm{nH}}(k)$. How to generalize the discussions to the frequency dependent cases, like $H_{S,{\rm eff}}(k,\omega)=H_{S}(k)+\Sigma_{S}(k,\omega)$? As far as we know, this problem has not yet been thoroughly explored in the literature. This leads to the second motivation of this work, that is, providing a comprehensive discussion on the complex frequency detections of the exact sub-system effective Hamiltonian, i.e., $H_{S,{\rm eff}}(k,\omega)$.
Starting from a exactly solvable model, our calculations reveal the following conclusion: while CFE, CFS, and CFF yield consistent outcomes under $H_{S,{\rm nH}}(k)$, they diverge when applied to $H_{S,{\rm eff}}(k,\omega)$. Specifically, as shown in Fig. 1 (c), the CFE and CFS yield the exact results shown in the right panels (with no non-Bloch response), whereas the CFF detects the approximated ones shown in the left panels (with the presence of non-Bloch response). Our work not only establishes a foundation for understanding the NHSE and non-Bloch dynamics in quantum systems but also provides critical guidance for experimental verification.
Model and subsystem Green’s function—Our Hamiltonian adopts the following Hermitian form:
$\hat{H}_{tot}=\hat{H}_{S}+\hat{H}_{S^{\prime}}+\hat{H}_{SS^{\prime}},$ (1)
where the hopping parameters are illustrated in Fig. 2 (a). The subsystem $S$ corresponds to a one-dimensional spinless Rice-Mele model with broken time-reversal symmetry [32], positioned in the $y=1$ region, with sublattices $A$ and $B$ represented by red and blue points, respectively. Applying Fourier transformation, we obtain the Bloch Hamiltonian:
$H_{S}(k)=(t_{1}+t_{2}\mathrm{cos}~{}k)\sigma_{x}+t_{2}\mathrm{sin}~{}k\sigma_{y}+(\lambda\mathrm{sin}~{}k-\mu_{S})\sigma_{z},$ (2)
where the Pauli matrix acts on the sublattice degree of freedom.
The subsystem $S^{\prime}$ comprises a two-dimensional single-band model inhabiting the $y\geq 2$ region, featuring anisotropic hoppings with $t_{x}\neq t_{y}$. Its $x$-direction Bloch Hamiltonian adopts the layer-resolved form:
$H_{S^{\prime}}(k)=\begin{pmatrix}h_{S^{\prime}}(k)&t_{y}\sigma_{0}&0&0&0&...\\ t_{y}\sigma_{0}&h_{S^{\prime}}(k)&t_{y}\sigma_{0}&0&0&...\\ 0&t_{y}\sigma_{0}&h_{S^{\prime}}(k)&t_{y}\sigma_{0}&0&...\\ ...&...&...&...&...&...\end{pmatrix},$ (3)
where $h_{S^{\prime}}(k)=(t_{x}+t_{x}\mathrm{cos}~{}k)\sigma_{x}+t_{x}\mathrm{sin}~{}k\sigma_{y}-\mu_{S^{\prime}}\sigma_{0}$ describes the Hamiltonian at each $y$-layer using a two-site unit cell. The interlayer hopping in the subsystem $S^{\prime}$ is parameterized by $t_{y}\sigma_{0}$, and the total number of $y$-layers in the subsystem $S^{\prime}$ is denoted as $N_{y}$.
For conceptual clarity, we restrict inter-subsystem coupling to the $A$-sublattice (denoted as $t_{A}$). In this case, as derived in the Supplementary Material (SM), the exact self-energy for the subsystem $S$ becomes:
$\Sigma_{S}(k,\omega)=\begin{pmatrix}\Sigma_{S}^{AA}(k,\omega)&0\\ 0&0\\ \end{pmatrix},$ (4)
where $\Sigma_{S}^{AA}(k,\omega)=t^{2}_{A}[1/(\omega-H_{S^{\prime}}(k))]_{11}$. Fig. 2 (b) and (c) present the corresponding density plots of the real and imaginary components of $\Sigma^{AA}_{S}(k,\omega)$, respectively.
Figure 3: (a) and (b) The spectral functions with and without NHA, respectively. (c) A comparison between the real frequency DOS with and without the NHA. (d) and (e) The complex frequency DOS with and without NHA, respectively, where the black circles indicate the poles of the CFGF. The parameters are identical to that in Fig. 2.
Based on this exact self-energy, the exact GF for subsystem $S$ under PBC can be written as:
$G_{S,{\rm eff}}^{R}(k,\omega)=\frac{1}{\omega-H_{S,\rm eff}(k,\omega)},$ (5)
where the frequency-dependent non-Hermitian Hamiltonian is explicitly expressed as
$H_{S,\rm eff}(k,\omega)=H_{S}(k)+\Sigma_{S}(k,\omega+i\gamma)-i\gamma.$ (6)
Here, the parameter $i\gamma$ follows the conventional retarded GF formalism, while $\omega$ spans the entire complex energy plane. In practical numerical calculations, $\gamma$ is chosen as a small value, which does not affect the qualitative results.
Non-Hermitian approximation—We now check the validity of the NHA. As shown in Fig. 2 (b) and (c), the energy band of $H_{S}(k)$ is plotted against the real and imaginary components of the self-energy, respectively. Notably, within the band dispersion region, the self-energy can be well approximated by a constant, i.e.,
$\Sigma_{S}^{AA}(k,\omega+i\gamma)\approx\Sigma_{S}^{AA}(0,i\gamma)+\mathcal{O}(k)+\mathcal{O}(\omega).$ (7)
This NHA yields the following approximated GF,
$G_{S,{\rm nH}}^{R}(k,\omega)=\frac{1}{\omega-H_{S,\rm nH}(k)},$ (8)
where the frequency-independent non-Hermitian Hamiltonian is given by:
$H_{S,\rm nH}(k)=H_{S}(k)+\begin{pmatrix}\Sigma_{S}^{AA}(0,i\gamma)&0\\ 0&0\\ \end{pmatrix}-i\gamma.$ (9)
To further verify the NHA, we present a comparative analysis of exact versus approximated spectral functions, i.e., $A_{S,\rm eff/nH}(k,\omega)=-(1/\pi)\mathrm{Im}~{}\mathrm{Tr}~{}G^{R}_{S,\rm eff/nH}(k,\omega\in\mathbb{R})$. As illustrated in Fig. 3 (a) and (b), the remarkable agreement between these spectral functions confirms the validity of the NHA. This conclusion is further corroborated by the density of states (DOS) at real frequencies shown in Fig. 3 (c), which follows the conventional definition, i.e., $\rho_{S,\rm{eff}/\rm{nH}}(\omega)=(1/N_{S})\sum_{k}A_{S,\rm{eff}/\rm{nH}}(k,\omega)$.
However, significant discrepancies emerge when considering the complex frequency domain. Fig. 1 (c) demonstrates a representative example of these differences. For comprehensive analysis, we introduce the complex-frequency DOS [100] across the entire complex plane, i.e.,
$D_{S,\rm eff/nH}(\omega)=-\frac{1}{N_{S}\pi}\sum\limits_{k}\mathrm{Im}~{}\mathrm{Tr}~{}G^{R}_{S,\rm eff/nH}(k,\omega\in\mathbb{C}),$ (10)
which can reveal the pole structures of the CFGF [100, 103]. As shown in Fig. 3 (d)-(e), the approximated GF displays two closed-loop poles, contrasting with the exact case that features a branch cut positioned at $\mathrm{Im}~{}\omega=-\gamma$. Due to the vanishing of the spectral winding in the exact poles under PBC, it is expected that there is no non-Bloch response in the GF, which explains the paradox observed in Fig. 1 (c).
In summary, although the NHA demonstrates excellent performance in simulating real-frequency GF, it fundamentally modifies the global pole structures. Therefore, it is necessary to examine the nontrivial roles of the NHA in variant complex-frequency detection methodologies.
Figure 4: (a) The poles of the OBC subsystem GF with and without NHA. (b) A comparison between the time-dependent response function for the CFE and the CFGF with and without NHA. Here $\omega_{7}=-1-0.05i$, $\omega_{8}=-1-0.115i$, $j_{0}=30$, and $i_{0}=36$. (c) A comparison between the CFE and CFS at $\omega_{7}$ with different synthesis domains and spectral resolution. Here $i_{0}=36$. (d) and (e) The complex frequency DOS for the CFF at $\omega_{r}=4$ and $\omega_{r}=5$, respectively. Other parameters are identical to that in Fig. 2.
Complex frequency detection—We now focus on the complex frequency detection of the subsystem. The theoretical framework originates from the following driving dissipative equation [100]:
$i\frac{d\langle\hat{\bm{a}}(t)\rangle}{dt}=(H_{tot}-i\gamma)\langle\hat{\bm{a}}(t)\rangle+\bm{F}(t),$ (11)
where $H_{tot}$ represents the first quantized Hamiltonian of $\hat{H}_{tot}$ in real space. The parameter $\gamma$ characterizes the uniform dissipation within the GF. $\langle\hat{\bm{a}}(t)\rangle=\{\langle\hat{\bm{a}}_{S}(t)\rangle,\langle\hat{\bm{a}}_{S^{\prime}}(t)\rangle\}^{T}$ describes the mean value of single-particle operators at each lattice site of the total system, and $\bm{F}(t)$ denotes the external drive. Experimentally, when $\bm{F}(t)$ is applied to the subsystem $S$, the corresponding response $\langle\hat{\bm{a}}_{S}(t)\rangle$ is detected. For example, consider an external driving defined as
$\bm{F}(t)=\theta(t)e^{-i\omega_{r}t}\{0,..,0,F_{S,i},0,...,0\}^{T},~{}~{}~{}~{}\omega_{r}\in\mathbb{R},$ (12)
which acts exclusively on the $i$th-site of the subsystem $S$. Following SM derivations, the induced response at the $j$-th site, i.e., $\langle\hat{a}_{S,j}(t)\rangle$, demonstrates proportionality to the real-frequency Green’s function in the long-time limit. This relationship can be expressed as:
$\lim_{t\rightarrow\infty}\langle\hat{a}_{S,j}(t)\rangle=\left[G_{S,{\rm eff}}^{R}(\omega_{r})\right]_{ji}F_{S,i}e^{-i\omega_{r}t},$ (13)
where $G_{S,\rm eff}^{R}(\omega_{r})$ denotes the exact GF of the subsystem defined in Eq. 5 (real space representation), and $\omega_{r}$ corresponds to the driving frequency. Given the predetermined values of $F_{S,i}$ and $\omega_{r}$, the real-frequency Green’s function becomes experimentally accessible through $\langle\hat{a}_{S,j}(t)\rangle$ detection. This methodology maintains alignment with classical wave system protocols for real-frequency Green’s function measurement [115, 116, 117, 118, 119].
Complex frequency excitation—When extending the driving frequency from real ($\omega_{r}$) to complex values ($\omega_{c}$) in Eq. 12, the concept of CFE emerges. As demonstrated in the SM, the solution to Eq. 11 under complex frequency driving
$\bm{F}(t)=\theta(t)e^{-i\omega_{c}t}\{0,..,0,F_{S,i},0,...,0\}^{T},~{}~{}~{}~{}\omega_{c}\in\mathbb{C},$ (14)
takes the form
$\langle\hat{a}_{S,j}(t)\rangle=\left[\chi_{S}(\omega_{c})\right]_{ji}F_{S,i}e^{-i\omega_{c}t},$ (15)
where the response function $\chi_{S}(\omega_{c})$ is expressed as
$\chi_{S}(\omega_{c})=$ $G^{R}_{S,\rm eff}(\omega_{c})-[G^{R}_{tot}(\omega_{c})e^{-i(H_{tot}-i\gamma-\omega_{c})t}]_{S}.$ (16)
Here, the subscript $S$ indicates the subsystem component, and $G^{R}_{tot}(\omega_{c})$ denotes the CFGF for the total system. We note that this solution assumes the initial condition $\langle\hat{\bm{a}}(0)\rangle=0$, whose value will not affect the subsequent conclusions, as proved in the SM.
A critical observation arises from the above formulation: when the complex frequency $\omega_{c}$ resides above the branch cut determined by $i\gamma$ (light pink region in Fig. 4 (a)), the corresponding response function will converge to the exact CFGF of the subsystem in the $t\rightarrow\infty$ limit, i.e.,
$\lim_{t\rightarrow\infty}\chi_{S}(\omega_{c})=G_{S,\rm eff}^{R}(\omega_{c}),~{}~{}~{}~{}\operatorname{Im}\omega_{c}>-\gamma.$ (17)
Conversely, frequencies below the $i\gamma$ line will yield divergent responses due to the second term’s divergence. As numerically verified in Fig. 4 (b), only the $\omega_{7}$ case converges to the exact CFGF (blue dashed lines).
For comparative analysis, we show the NHA-predicted CFGF as red dashed lines, which exhibit clear deviations from our calculated response functions. This discrepancy stems from the modification of GF pole structures under the NHA (Fig. 3 (d)-(e)), rendering it invalid for frequencies distant from the real axis. Fig. 4 (b) explicitly demonstrates this limitation: for $\omega_{7}$, converged results deviate from NHA predictions; $\omega_{8}$ exhibits response divergence despite NHA suggesting convergence.
Key conclusions are summarized in Table 1: (i) The CFE exclusively detects the exact CFGF, i.e., Eq. 5, above the total Hamiltonian’s imaginary gap (light pink region); (ii) NHA fails in regions far from the real axis; (iii) Absence of non-Bloch response in long-time dynamics due to vanishing of spectral winding number.
Table 1: Results of CFE, CFS, and CFF for $t\rightarrow\infty$
Effective Region CFGF NHA Non-Bloch Resonse
[…]
CFS the same as CFE
[…]
    $\Gamma$ characterizes the imaginary gap. In the exact case, $\Gamma=-\gamma$.
Complex frequency synthesis—We now examine the CFS, a method based on the following observation: within the framework of CFE, the nonzero component of $\bm{F}(t)$ in Eq. 14 can be transformed through Fourier analysis as
[…]
where $\mathrm{Im}~{}\omega_{c}<0$ and $t>0$. Experimentally measured real frequency responses enable comprehensive reconstruction of the CFE solution through signal synthesis:
$\langle\hat{\bm{a}}_{S,\omega_{c}}(t)\rangle=\frac{i}{2\pi}\sum\limits^{\omega_{b}}_{\omega=\omega_{a}}\frac{\Delta\omega}{\omega-\omega_{c}}\langle\hat{\bm{a}}_{S,\omega}(t)\rangle,$ (19)
where $[\omega_{a},\omega_{b}]$ defines the synthesis domain and $\Delta\omega$ specifies the spectral resolution.
As summarized in Tab. 1, CFS demonstrates theoretical equivalence with CFE. A quantitative validation is presented in Fig. 4 (c), where the result derived from CFE is compared with synthesized responses across progressively expanding synthesis domains $\Delta\omega$ and refined spectral resolutions. The convergence toward the CFE result confirms the validity of the CFS method.
Complex frequency fingerprint—To capture the non-Bloch response, the recently developed CFF method under real-frequency driving is necessary. As demonstrated in the SM, as $t\rightarrow\infty$, the CFF specifically measures the following double-frequency GF,
$G^{\rm CFF}_{S,\rm eff}(k,\omega_{c},\omega_{r})=\frac{1}{\omega_{c}-H_{S,\rm eff}(k,\omega_{r})},$ (20)
where $\omega_{c}\in\mathbb{C}$ represents a complex parameter and $\omega_{r}\in\mathbb{R}$ corresponds to the driving frequency defined in Eq. 12. Crucially, for any fixed driving parameter $\omega_{r}$, this double-frequency GF precisely corresponds to the CFGF associated with the non-Hermitian Hamiltonian $H_{S,\rm eff}(k,\omega_{r})$. As $\omega_{r}$ evolves, the corresponding double-frequency GF changes accordingly. Fig. 4 (d) and (e) exemplify this behavior through the following complex-frequency DOS
$D_{S,\rm eff}^{\mathrm{CFF}}(\omega_{c},\omega_{r})=-\frac{1}{N_{S}\pi}\sum_{k}\mathrm{Im}~{}\mathrm{Tr}~{}G^{\mathrm{CFF}}_{S,\rm{eff}}(k,\omega_{c},\omega_{r})$ (21)
at frequencies $\omega_{r}=4$ and $\omega_{r}=5$, respectively.
For any given $\omega_{r}$, once $H_{S,\rm eff}(k,\omega_{r})$ exhibits the NHSE, the corresponding non-Bloch response can be detected via the double frequency GF by setting $\omega_{c}$ with a nonzero spectral winding. An example is illustrated on the left side of Fig. 1 (c) with $\omega_{r}=0$.
In summary, as illustrated in Tab. 1: (i) The CFF methodology effectively identifies double-frequency GF governed by Eq. 20; (ii) For each driving frequency $\omega_{r}$, the CFF is described by the non-Hermitian Hamiltonian at $\omega_{r}$, i.e., $H_{S,\rm eff}(k,\omega_{r})$; (iii) This framework establishes a unified platform for the systematic exploration of NHSE and non-Bloch responses in a Hermitian quantum system.
Conclusion.—In this work, we have established a framework for detecting complex frequency responses in a frequency-dependent Hamiltonian of the subsystem. Our comparative study reveals that while both CFE and CFS methods yield identical results characterizing the exact CFGF without NHSE signatures, the CFF approach uniquely manifests non-Bloch responses through its detection of double-frequency GFs. Our work elucidates the physical origin of non-Hermitian Hamiltonians in quantum systems through exact theoretical formalisms, bypassing traditional approximation constraints. These advancements provide foundational insights for decoding non-Hermitian phenomena in quantum systems and offer concrete guidance for experimental implementations.
J. Huang and Z. Yang were sponsored by the National Key R&\&&D Program of China (No. 2023YFA1407500) and the National Natural Science Foundation of China (No. 12322405, 12104450, 12047503). J. Hu was sponsored by the Ministry of Science and Technology (Grant No. 2022YFA1403901), National Natural Science Foundation of China ( No. 12494594), and the New Cornerstone Investigator Program.